OpenStudy (anonymous):

is this true for a given even function:

OpenStudy (anonymous):

\[\int\limits_{-a}^{a}f(x) dx = 2 \int\limits_{0}^{a} f(x) dx\]

OpenStudy (anonymous):

:) it is

OpenStudy (anonymous):

so, if integral of even = odd lets assume g(x) = integral of f(x) we get this: \[g(a) - g(-a) = 2g(a) - 2g(0)\] and since g is odd \[2g(a)= 2g(a) -2g(0)\] does this mean g(0) must always be zero ? ie g always passes through the origin?

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

if the integral is odd .. it's zero if even 2 times _0^a I

OpenStudy (anonymous):

integral has to be odd if f(x) is even

OpenStudy (experimentx):

there is a sign change.

OpenStudy (anonymous):

where?

OpenStudy (experimentx):

can you tell me what this means? "integral of even = odd"?

OpenStudy (experimentx):

this goes like this ... we split up the integral \[ \int_{-a}^0f(x) dx - \int_0^af(x) dx\]

OpenStudy (anonymous):

if f(-x) = f(x) then g(-x) = -g(x)

OpenStudy (experimentx):

hold on ... i guess i get what you mean.

OpenStudy (experimentx):

guess that's wrong.

OpenStudy (experimentx):

it seems for x^2 ... the integration yields x^3 which is odd same goes for cos(x)

OpenStudy (anonymous):

well it's established that integrating an even gets an odd and vice versa, try and find the integral of an even function, which does not pass through the origin like integrating cos(x) gets sin(x) , and sin passes through the origin. my conjecture is that there is no function f(x) such that its integral g(x) does not pass through the origin. i dont know whether my first argument is valid.

OpenStudy (experimentx):

let's try with abs int abs(x) dx = x^2/2 if x>0 both cases ... but this is a piece wise function.

OpenStudy (anonymous):

i have to go to work, be back later probably

OpenStudy (anonymous):

bye!

OpenStudy (experimentx):

i'll try to think and post.

OpenStudy (experimentx):

ahh.. .man too sleepy can't think much :( sorry